On Sunday 845 people went to the Zoo. On Monday only 169 people went. What is the per cent decrease in the people visiting the zoo on Monday

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Given that on Sunday, 845 people went to the zoo and on Monday, 169 people went.
Decrease in the numberof people = 845 − 169 = 676
Percentage decrease =(Decrease in the number of people ×100) %
Number of people who went to zoo on sunday
= (676 ×100) %
845
= 80 %
Thanks